OpenStudy (anonymous):

find the value of x . a) x=18 b) x=22.5 c) x=30 d) x=54

pooja195 (pooja195):

Where is the equation?

pooja195 (pooja195):

Vertical angles are equal right? So just set x+90=4x <----from here solve for x :)

pooja195 (pooja195):

Understood?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well no

OpenStudy (anonymous):

is it b

pooja195 (pooja195):

Yes :)

pooja195 (pooja195):

Do you understand why?

OpenStudy (anonymous):

yes because if u divide 90 by 4 u get 45/2 then u get 22.5

pooja195 (pooja195):

o.O nope

pooja195 (pooja195):

You have the right answer but you don't know why that is the answer

OpenStudy (anonymous):

no

pooja195 (pooja195):

x+90=4x <----from here solve for x :) all you do is subtract x from both sides 4x-1x=3 right? then all you do is divide 90=3x divide both sides by 3 90/3 = 30 3/3 is 1 so just x therefore x=30 understood? Or need help?

OpenStudy (anonymous):

yes i understand

pooja195 (pooja195):

good :)